# Physics-Guided Multi-Fidelity DeepONet for Data-Efficient Flow Field Prediction

### v1

@@ Abstract
This study presents an enhanced multi-fidelity deep operator network (DeepONet) framework for efficient spatio-temporal flow field prediction, with particular emphasis on practical scenarios where high-fidelity data is scarce. We introduce several key innovations to improve the framework’s efficiency and accuracy. First, we enhance the DeepONet architecture by incorporating a merge network that enables more complex feature interactions between operator and coordinate spaces, achieving a 50.4% reduction in prediction error compared to traditional dot-product operations. We further optimize the architecture through temporal positional encoding and point-based sampling strategies, achieving a 7.57% improvement in prediction accuracy while reducing training time by 96% through efficient sampling and automatic mixed precision training. Building upon this foundation, we develop a transfer learning-based multi-fidelity framework that leverages knowledge from pre-trained low-fidelity models to guide high-fidelity predictions. Our approach freezes the pre-trained branch and trunk networks while making only the merge network trainable during high-fidelity training, preserving valuable low-fidelity representations while efficiently adapting to high-fidelity features. Through systematic investigation, we demonstrate that this fine-tuning strategy not only significantly outperforms linear probing and full-tuning alternatives but also surpasses conventional multi-fidelity frameworks by up to 76%, while achieving up to 43.7% improvement in prediction accuracy compared to single-fidelity training. The core contribution lies in our novel time-derivative guided sampling approach, which strategically selects high-fidelity training points based on temporal dynamics identified by pre-trained low-fidelity models. This physics-guided sampling strategy demonstrates remarkable effectiveness: it maintains prediction accuracy equivalent to models trained with the full dataset while requiring only 60% of the original high-fidelity samples. Through comprehensive experiments across multiple resolutions and dataset sizes, we show that our enhanced framework offers a practical solution for industrial applications where high-fidelity data acquisition is costly and time-consuming.
Keywords Deep operator networks  $\cdot$ Multi-fidelity modeling  $\cdot$ Transfer learning  $\cdot$ Time-derivative guided sampling  $\cdot$ Small dataset  $\cdot$ Spatio-temporal flow prediction
@@ 1 Introduction
In the era of scientific machine learning (SciML), accurate and efficient prediction of spatio-temporal flow fields plays a crucial role across various engineering disciplines, particularly in the development of digital twins that require real-time flow field predictions for virtual replicas of physical systems. While traditional numerical approaches like computational fluid dynamics (CFD) offer high accuracy, their computational demands make them impractical for real-time applications, where rapid prediction and flexible adaptation to different flow parameters are essential. This has led to the emergence of various deep learning approaches, including multilayer perceptrons (MLPs) [30, 35], convolutional neural networks (CNNs) [10, 19, 7, 5, 9, 4], and graph neural networks (GNNs) [23, 38, 12, 37], as surrogate models for flow field prediction. However, these architectures face inherent limitations: MLPs lack spatial structure awareness and cannot flexibly handle varying mesh resolutions, requiring retraining for each new mesh configuration—a significant drawback for digital twins that must adapt to different system configurations. CNNs are constrained by fixed grid resolutions, limiting their applicability to irregular domains. GNNs, while capable of handling unstructured meshes, require extensive memory resources to store and process mesh connectivity information, making them computationally intensive for large-scale problems [22].
@@
The deep operator network (DeepONet) [16], introduced to learn mappings between function spaces, offers a compelling alternative for spatio-temporal flow field prediction tasks by overcoming the limitations of traditional methods [8, 1, 17, 26, 25]. Its unique architecture, consisting of separate branch and trunk networks, enables prediction at arbitrary spatial locations without being constrained by specific grid resolutions or mesh connectivities. This resolution-invariant property makes DeepONet particularly suitable for flow field predictions where data may come from simulations with varying mesh configurations. However, the standard DeepONet architecture may not be ideally designed to capture the complex dynamics and intricate features of spatio-temporal behavior [1, 13, 11, 15]. To fully exploit DeepONet’s potential for spatio-temporal flow field prediction, we recognize the need for specialized architectural refinements and training methodologies tailored to fluid flow applications. By systematically developing and integrating enhancements to the core DeepONet framework, we aim to significantly improve both the accuracy and computational efficiency of flow field predictions through: 1) expanding the architecture with additional networks for more flexible feature interactions between operator and coordinate spaces, 2) addressing spectral bias through strategic positional encoding implementation, and 3) implementing efficient point-based sampling techniques with automatic mixed precision training to reduce computational overhead.
@@
However, training DeepONet remains challenging, especially when high-fidelity data is limited due to computational constraints. In this context, multi-fidelity modeling has emerged as a powerful approach in SciML, effectively leveraging abundant low-fidelity data alongside limited high-fidelity data to enhance prediction accuracy while maintaining computational efficiency [6, 20, 36, 39, 34, 33]. Existing multi-fidelity DeepONet frameworks [18, 2, 3] predominantly employ coupled architectures, where low-fidelity (LF) and high-fidelity (HF) networks are trained and utilized together. This coupling approach fundamentally requires identical query points for both branch inputs (function queries) and trunk inputs (spatio-temporal coordinates) across different fidelity levels. This severely restricts DeepONet’s primary advantage: the flexibility to leverage operators at arbitrary points. It forces researchers to design low-fidelity and high-fidelity experiments with identical branch inputs and trunk inputs beforehand—a requirement that is highly inefficient and often impractical in industrial applications where data collection follows operational constraints rather than organized experimental design. This coupled architecture introduces significant computational limitations, requiring sequential two-network inference processes that increase memory requirements and computational costs. The inefficiency scales poorly with multiple fidelity levels, as each additional fidelity layer requires its own network, creating cascading overhead particularly problematic when integrated with physics-informed methods, where automatic differentiation should propagate through all architectures.
@@
Beyond these architectural limitations, fundamental methodological challenges exist in effectively leveraging the relationship between LF and HF data during DeepONet training. While Xu et al. [32] proposed a two-phase approach pre-training a LF DeepONet before training a residual DeepONet for high-fidelity predictions—which removes the requirement of identical LF and HF datasets due to its sequential transfer learning approach—their framework still necessitates two separate networks during both training and inference. This decoupling of training between LF and HF DeepONets allows for more flexibility in dataset design, but continues to suffer from computational inefficiency as both models must be employed during inference. Additionally, it lacks a systematic mechanism to identify critical regions where high-fidelity training would be most beneficial, resulting in potentially inefficient use of the pre-trained low-fidelity model and valuable high-fidelity data. Combined with the inherent drawbacks of residual-based architectures discussed earlier, these issues highlight the need for a fundamentally different approach to multi-fidelity operator learning.
A more refined approach to multi-fidelity DeepONet for spatio-temporal flow field prediction would incorporate four key innovations: (1) a unified network structure that eliminates the need for coupled use of low-fidelity and high-fidelity DeepONets during inference, reducing computational overhead and memory requirements; (2) preservation of DeepONet’s inherent query flexibility by removing the requirement for low-fidelity and high-fidelity datasets to share identical function inputs or spatio-temporal query points; (3) a physics-guided sampling strategy that identifies regions of significant flow dynamics using pre-trained low-fidelity models, strategically concentrating high-fidelity data collection in these dynamically critical areas to maximize information gain while minimizing computational cost; and (4) comprehensive evaluation of the relationship between dataset sizes and multi-fidelity model performance, with particular emphasis on strategies for minimizing high-fidelity data requirements while maintaining prediction accuracy.
@@
    We propose an expanded DeepONet architecture incorporating a merge network for complex feature interactions between operator (branch net) and coordinate spaces (trunk net). Through extensive experiments, we demonstrate that our element-wise multiplication merge strategy achieves a 50.4% reduction in prediction error compared to traditional dot-product operations, while maintaining similar model parameters. This significant improvement highlights the importance of non-linear feature processing in capturing intricate flow dynamics.
@@
    We address spectral bias in DeepONet’s coordinate processing through strategic positional encoding implementation. Our comprehensive analysis reveals that selective temporal coordinate encoding achieves a 7.57% improvement in prediction accuracy, while applying encoding to all spatio-temporal coordinates can lead to instability. This investigation provides crucial insights into optimizing coordinate representation while maintaining training stability.
@@
    We leverage DeepONet’s unique point-based prediction capability to develop efficient training protocols. Our approach demonstrates that training with a subset of spatial points significantly reduces computational overhead while maintaining prediction accuracy—challenging the assumption that more training points always lead to better results. Combined with automatic mixed precision training, we achieve a remarkable 96% reduction in training time (from 8.672 to 0.350 hours) while maintaining or even enhancing prediction accuracy, establishing new benchmarks for practical DeepONet implementation.
@@
    We introduce a transfer learning-based MF-DeepONet framework that fundamentally departs from sequential, two-network approaches like those in [18, 2, 3]. By strategically transferring knowledge from pre-trained low-fidelity models to a single unified network for high-fidelity predictions, our approach completely eliminates both the computational overhead of maintaining coupled networks and the restrictive constraint of matching query points between fidelity levels. This critical advancement preserves DeepONet’s fundamental capability to evaluate operators at arbitrary points while removing the need for aligned branch and trunk inputs across fidelity levels—enabling seamless integration of heterogeneous data from different sources and resolutions. In direct comparison with conventional multi-fidelity approaches proposed by Lu et al. [18], our framework demonstrates dramatically superior performance with up to 76% improvement in prediction accuracy while significantly reducing computational requirements during both training and inference.
@@
    Through systematic investigation of different transfer learning configurations—including fine-tuning (where only the merge network is trainable while branch and trunk networks are frozen), full-tuning (where all network components are retrained), and linear probing (where only the final layer is trainable)—we demonstrate that the selective fine-tuning approach significantly outperforms alternatives. Our extensive experimentation with various low-fidelity resolutions (16×16, 32×32, 64×64) and dataset sizes (50-300 samples) reveals that fine-tuning achieves up to 43.7% improvement in prediction accuracy compared to single-fidelity training, while maintaining computational efficiency by preserving valuable representations learned from low-fidelity data.
@@
    We propose a physics-guided sampling strategy based on temporal derivatives that efficiently identifies critical regions requiring high-fidelity data. This intuitive approach demonstrates remarkable effectiveness: our method reduces prediction error by 10.02-20.73% compared to traditional uniform sampling methods using the same amount of data. Furthermore, using only 60% of high-fidelity samples with the same low-fidelity data, we achieve similar prediction accuracy compared to models trained with the full high-fidelity dataset without sampling. These findings represent a significant breakthrough for industrial applications where high-fidelity data acquisition is prohibitively expensive or time-consuming, providing a principled methodology for dramatically reducing computational requirements without sacrificing prediction accuracy.
The remainder of this paper is organized as follows. Section 2 presents the theoretical foundations of our enhanced DeepONet architecture and multi-fidelity framework, including the merge network design and time-derivative guided sampling strategy. Section 3 describes the flow field data generation process with varying initial conditions and fidelity levels. Section 4 evaluates the architectural enhancements to DeepONet, analyzing the effects of different merge processes, positional encoding, and efficient training techniques. Section 5 investigates our MF-DeepONet framework, comparing different transfer learning approaches and demonstrating performance with limited high-fidelity data. Section 6 explores the time-derivative guided sampling strategy and quantifies its effectiveness in reducing high-fidelity data requirements. Finally, Section 7 summarizes our findings and discusses implications for future research.
@@ 2.1.1 Enhanced DeepONet with Merge Network
The original DeepONet consists of two main components: a branch network and a trunk network. The branch network maps the input function information (e.g., initial conditions, boundary conditions, or other function characterizations) to a latent representation $\mathbf{B}(u)\in\mathbb{R}^{p}$, where $p$ is the dimension of the latent space. The trunk network takes spatio-temporal coordinates, $(x,y,t)$ in this study, as input and outputs latent space $\mathbf{T}(y)\in\mathbb{R}^{p}$. In the standard DeepONet formulation, these two latent representations are combined through a simple dot product operation to produce the final prediction: $f(u)(y)=\mathbf{B}(u)\cdot\mathbf{T}(y)$.
@@
    Final $\text{Final output}=\mathbf{B}(u)\cdot\mathbf{T}(y)$ (1)
@@
    Final $\text{Final output}=\mathbf{B}(u)\cdot\mathbf{T}(y)+b$ (2)
@@
    Final $\text{Final output}=\text{MergeNet}(\mathbf{B}(u)\odot\mathbf{T}(y))$ (3)
@@
    Final $\text{Final output}=\text{MergeNet}([\mathbf{B}(u);\mathbf{T}(y)])$ (4)

### v2

# Data-Efficient Deep Operator Network for Unsteady Flow: A Multi-Fidelity Approach with Physics-Guided Subsampling

@@ Abstract
This study presents an enhanced multi-fidelity deep operator network (DeepONet) framework for efficient spatio-temporal flow field prediction, with particular emphasis on practical scenarios where high-fidelity data is scarce. We introduce several key innovations to improve the framework’s efficiency and accuracy. First, we enhance the DeepONet architecture by incorporating a merge network that enables more complex feature interactions between operator and spatio-temporal coordinates, achieving a 50.4% reduction in prediction error compared to traditional dot-product operations. We further optimize the architecture through temporal positional encoding and point-based subsampling strategies, achieving a 7.57% improvement in prediction accuracy while reducing training time by 96%. Building upon this foundation, we develop a transfer learning-based multi-fidelity framework that leverages knowledge from pre-trained low-fidelity models to guide high-fidelity predictions. Our approach freezes the pre-trained branch and trunk networks while making only the merge network trainable during high-fidelity training, preserving valuable low-fidelity representations while efficiently adapting to high-fidelity features. Through systematic investigation, we demonstrate that this fine-tuning strategy not only outperforms linear probing and full-tuning alternatives but also surpasses conventional multi-fidelity frameworks by up to 76%, while achieving up to 43.7% improvement in prediction accuracy compared to single-fidelity training. As another core contribution, we introduce a physics-guided subsampling approach, which strategically selects high-fidelity training spatial points based on temporal dynamics identified by pre-trained low-fidelity models. This physics-guided subsampling strategy demonstrates remarkable effectiveness, achieving prediction accuracy comparable to a baseline model while reducing the high-fidelity sample requirement by 40%. The robustness of these contributions, particularly our proposed MF-DeepONet framework and time-derivative subsampling, is further validated on a second, more complex dataset where they continue to demonstrate superior or highly competitive performance against conventional benchmarks. Through comprehensive experiments across multiple resolutions and two distinct datasets of increasing complexity, our enhanced framework demonstrates the potential to reduce the required high-fidelity dataset size while maintaining predictive accuracy.
Keywords Deep operator networks  $\cdot$ Multi-fidelity modeling  $\cdot$ Transfer learning  $\cdot$ Physics-guided subsampling  $\cdot$ Small dataset  $\cdot$ Spatio-temporal flow prediction
@@ 1 Introduction
In the era of scientific machine learning (SciML), accurate and efficient prediction of spatio-temporal flow fields plays a crucial role across various engineering disciplines, particularly in the development of digital twins that require real-time flow field predictions for virtual replicas of physical systems. While traditional numerical approaches like computational fluid dynamics (CFD) offer high accuracy, their computational demands make them impractical for real-time applications, where rapid prediction and flexible adaptation to different flow parameters are essential. This has led to the emergence of various deep learning approaches, including multilayer perceptrons (MLPs) [46, 52], convolutional neural networks (CNNs) [19, 30, 14, 12, 17, 11], and graph neural networks (GNNs) [36, 55, 21, 54], as surrogate models for flow field prediction. However, these architectures face inherent limitations: MLPs lack spatial structure awareness and cannot flexibly handle varying mesh resolutions, requiring retraining for each new mesh configuration—a significant drawback for digital twins that must adapt to different system configurations. CNNs are constrained by fixed regular grid resolutions, limiting their applicability to irregular domains. GNNs, while capable of handling unstructured meshes, require extensive memory resources to store and process mesh connectivity information, making them computationally intensive for large-scale problems [34].
@@
The deep operator network (DeepONet) [26], introduced to learn mappings between function spaces, offers a compelling alternative for spatio-temporal flow field prediction tasks by overcoming the limitations of traditional methods [15, 2, 28, 41, 38]. Its unique architecture, consisting of separate branch and trunk networks, enables prediction at arbitrary spatial locations without being constrained by specific grid resolutions or mesh connectivities. This resolution-invariant property makes DeepONet particularly suitable for flow field predictions where data may come from simulations with varying mesh configurations. However, the standard DeepONet architecture may not be ideally designed to capture the complex dynamics and intricate features of spatio-temporal behavior [2, 22, 20, 24]. To fully exploit DeepONet’s potential for spatio-temporal flow field prediction, we recognize the need for specialized architectural refinements and training methodologies tailored to fluid flow applications. By systematically developing and integrating enhancements to the core DeepONet framework, we aim to significantly improve both the accuracy and computational efficiency of flow field predictions through: 1) expanding the architecture with additional networks for more flexible feature interactions between operator and spatio-temporal coordinates, 2) addressing spectral bias through strategic positional encoding implementation, and 3) implementing efficient point-based subsampling techniques with automatic mixed precision training to reduce computational overhead.
@@
However, training DeepONet remains challenging, especially when high-fidelity data is limited due to computational constraints. In this context, multi-fidelity modeling has emerged as a powerful approach in SciML, effectively leveraging relatively abundant low-fidelity data alongside limited high-fidelity data to enhance prediction accuracy while maintaining computational efficiency [13, 32, 53, 56, 51, 50]. Existing multi-fidelity DeepONet frameworks [29, 9, 10] predominantly employ coupled architectures, where low-fidelity (LF) and high-fidelity (HF) networks are trained and utilized together. This coupling approach fundamentally requires identical query points for both branch inputs (function queries) and trunk inputs (spatio-temporal coordinates) across different fidelity levels. This severely restricts DeepONet’s primary advantage: the flexibility to leverage operators at arbitrary points. It forces researchers to design low-fidelity and high-fidelity experiments with identical branch inputs and trunk inputs beforehand—a requirement that is highly inefficient and often impractical in industrial applications where data collection follows operational constraints rather than organized experimental design. While the multi-fidelity DeepONet proposed by Howard et al. [16] successfully addressed this constraint by adopting a multi-fidelity deep neural network approach [31, 56], it still requires the coupled use of LF and HF networks during inference. This coupled architecture introduces computational overhead by requiring sequential two-network inference, which increases both memory requirements and computational costs. The inefficiency scales poorly with multiple fidelity levels, as each additional fidelity layer requires its own network, creating cascading overhead particularly problematic when integrated with physics-informed methods, where automatic differentiation should propagate through all architectures.
@@
Beyond these architectural limitations, fundamental methodological challenges exist in effectively leveraging the relationship between LF and HF data during DeepONet training. While Xu et al. [49] proposed a two-phase approach pre-training an LF DeepONet before training a residual DeepONet—which removes the requirement of identical LF and HF datasets by using its sequential transfer learning approach—their framework still necessitates two separate networks during inference stage. This decoupling between LF and HF DeepONets during training phase allows for more flexibility in dataset design, but continues to suffer from computational inefficiency as both models must be employed during inference. Additionally, it lacks a systematic mechanism to identify critical regions where high-fidelity training would be most beneficial, resulting in potentially inefficient use of the pre-trained low-fidelity model and valuable high-fidelity data. Combined with the inherent drawbacks of residual-based architectures discussed earlier, these issues highlight the need for a fundamentally different approach to multi-fidelity operator learning.
A more refined approach to multi-fidelity DeepONet for spatio-temporal flow field prediction would incorporate four key innovations: (1) a unified network structure that eliminates the need for coupled use of low-fidelity and high-fidelity DeepONets during inference, reducing computational overhead and memory requirements; (2) preservation of DeepONet’s inherent query flexibility by removing the requirement for low-fidelity and high-fidelity datasets to share identical function inputs or spatio-temporal query points; (3) a physics-guided subsampling strategy that identifies regions of significant flow dynamics using pre-trained low-fidelity models, strategically focusing the training process on high-fidelity data from these dynamically critical areas to maximize information gain while minimizing computational cost; and (4) comprehensive evaluation of the effects of dataset fidelities and sizes on multi-fidelity model performance, with particular emphasis on strategies for minimizing high-fidelity data requirements while maintaining prediction accuracy.
@@
    We propose an expanded DeepONet architecture incorporating a merge network for complex feature interactions between operator (branch net) and spatio-temporal coordinate spaces (trunk net). Through extensive experiments, we demonstrate that our element-wise multiplication merge strategy achieves a 50.4% reduction in prediction error compared to traditional dot-product operations, while maintaining similar model parameters. This significant improvement highlights the importance of non-linear feature processing in capturing intricate flow dynamics.
@@
    We systematically investigate positional encoding to establish practical guidelines for its application to spatio-temporal DeepONets. Our analysis reveals a crucial insight for physical systems: contrary to common practice in computer vision domains, encoding spatial coordinates degrades performance, while selective temporal encoding is beneficial, yielding a 7.57% improvement in prediction accuracy. This provides a valuable, non-obvious guideline for optimizing coordinate representation in spatio-temporal operator learning.
@@
    We leverage DeepONet’s unique point-based prediction capability to develop efficient training protocols. Our approach demonstrates that training with a subset of spatial points significantly reduces computational overhead while maintaining prediction accuracy—challenging the assumption that more training points can lead to better results. Combined with automatic mixed precision training, we achieve a remarkable 96% reduction in training time (from 8.672 to 0.350 hours) while maintaining or even enhancing prediction accuracy.
@@
    We introduce a transfer learning-based MF-DeepONet framework that fundamentally departs from sequential, two-network approaches like those in [29, 9, 10]. By strategically transferring knowledge from pre-trained low-fidelity models to a single unified network for high-fidelity predictions, our approach eliminates both the computational overhead of maintaining coupled networks and the restrictive constraint of matching query points between fidelity levels: where the transfer learning strategy is inspired by previous works that have successfully applied transfer learning techniques to multi-fidelity modeling [35, 39, 5, 25, 7, 18, 8, 1]. This critical advancement preserves DeepONet’s fundamental capability to evaluate operators at arbitrary points while removing the need for aligned branch and trunk inputs across fidelity levels—enabling seamless integration of heterogeneous data from different sources and resolutions. In direct comparison with conventional multi-fidelity approaches proposed by Lu et al. [29], our framework demonstrates superior performance with up to 76% improvement in prediction accuracy while significantly reducing computational requirements during both training and inference, and the robustness of such performance is further validated on a second, more challenging high-resolution dataset.
@@
    Through systematic investigation of different transfer learning configurations—including fine-tuning (where only the merge network is trainable while branch and trunk networks are frozen), full-tuning (where all network components are retrained), and linear probing (where only the final layer of the merge network is trainable)—we demonstrate that the selective fine-tuning approach significantly outperforms alternatives. Our extensive experimentation with various low-fidelity resolutions (16×16, 32×32, 64×64) and dataset sizes (50-300 samples) reveals that fine-tuning achieves up to 43.7% improvement in prediction accuracy compared to single-fidelity training, while maintaining computational efficiency.
@@
    We propose a physics-guided subsampling strategy based on temporal derivatives that efficiently identifies critical spatial areas requiring high-fidelity training. This intuitive approach demonstrates remarkable effectiveness: our method reduces prediction error by 10.02-20.73% compared to traditional uniform sampling methods using the same amount of data. Furthermore, using only 60% of high-fidelity samples with the same low-fidelity data, we achieve similar prediction accuracy compared to models trained without physics-guided sampling. The robustness of this physics-guided subsampling is also demonstrated on a more complex, higher-resolution dataset, where it remains an effective strategy that performs competitively against conventional residual-based sampling. These findings suggest a promising direction for reducing reliance on dense high-fidelity data in multi-fidelity modeling.
The remainder of this paper is organized as follows. Section 2 presents the theoretical foundations of our enhanced DeepONet architecture and multi-fidelity framework, including the merge network design and time-derivative guided subsampling strategy. Section 3 describes the flow field data generation process with varying initial conditions and fidelity levels. Section 4 evaluates the architectural enhancements to DeepONet, analyzing the effects of different merge processes, positional encoding, and efficient training techniques. Section 5 investigates our MF-DeepONet framework, comparing different transfer learning approaches and demonstrating performance with limited high-fidelity data. Section 6 explores the time-derivative guided subsampling strategy and quantifies its effectiveness in reducing high-fidelity data requirements. Section 7 further validates our framework on a second, more complex dataset and provides a comparative analysis of different guided sampling strategies. Finally, Section 8 summarizes our findings and discusses implications for future research.
@@ 2.1.1 Enhanced DeepONet with Merge Network
The original DeepONet consists of two main components: a branch network and a trunk network. The branch network maps the input function information (e.g., initial conditions, boundary conditions, or other function characterizations) to a latent representation $\mathbf{B}(\mathbf{u})\in\mathbb{R}^{p}$, where $\mathbf{u}$ is input function and $p$ is the dimension of the latent space. The trunk network takes spatio-temporal coordinates, $\mathbf{x}$; and it outputs latent space, $\mathbf{T}(\mathbf{x})\in\mathbb{R}^{p}$. In the standard DeepONet formulation, these two latent representations are combined through a simple dot product operation to produce the final prediction: $f(\mathbf{u})(\mathbf{x})=\mathbf{B}(\mathbf{u})\cdot\mathbf{T}(\mathbf{x})$ where function $f$ is an operator modeled by DeepONet.
@@
    Final $\text{Final output}=\mathbf{B}(\mathbf{u})\cdot\mathbf{T}(\mathbf{x})$ (1)
@@
    Final $\text{Final output}=\mathbf{B}(\mathbf{u})\cdot\mathbf{T}(\mathbf{x})+b$ (2)
@@
    Final $\text{Final output}=\text{MergeNet}(\mathbf{B}(\mathbf{u})\odot\mathbf{T}(\mathbf{x}))$ (3)
@@
    Final $\text{Final output}=\text{MergeNet}([\mathbf{B}(\mathbf{u});\mathbf{T}(\mathbf{x})])$ (4)